OpenStudy (shealyn):

I want to build a right triangular deck behind my house. The hypotenuse of the deck will be 25 feet long. The other two sides of the deck will be equal in length. What is the length of the deck?

OpenStudy (shealyn):

why

OpenStudy (shealyn):

haha

OpenStudy (steve816):

Well anyway back to mathematics

OpenStudy (shealyn):

so how do i find the length

OpenStudy (steve816):

use something called the pythagoreans theorem.

OpenStudy (steve816):

2x^2 = 25^2

OpenStudy (steve816):

solve for x

OpenStudy (mathmale):

if a and b represent the lengths of the legs (shorter sides) of a right triangle, and c represents the length of the hypo, then\[a^2 +b^2=c^2.\]

OpenStudy (shealyn):

i got 12.5 but thts not ann option

OpenStudy (mathmale):

this is the Pythagorean Theorem. If the 2 shorter sides of the triangle are equal, then a=b, and \[a^2+a^2=c^2.\] simplifiy this. Please show your work if you want feedback on it.

OpenStudy (shealyn):

woah if we even talk about it, it gets deleted 0O0

OpenStudy (mathmale):

Hint: c=25 feet. Yes, jokes that have no bearing on the solution of the problem at hand do get deleted, or the posters get warned. Let's focus on the problem solving, please.

OpenStudy (shealyn):

ik that c is 25 and c squared is 625 but how does that help if i dont have an a or b

OpenStudy (shealyn):

idk how to find a or b

OpenStudy (steve816):

"the other two sides of the deck will be equal in length" Which means a = b

OpenStudy (mathmale):

The whole point is to determine the length of one of the 2 shorter sides. Since these sides are the same length, finding a means you've also found b.

OpenStudy (mathmale):

this is a problem in algebra, where a and b represent unknowns. \[a^2+b^2=c^2\] but here a = b, and therefore,

OpenStudy (mathmale):

\[2a^2=625\]

OpenStudy (mathmale):

Solve this for a.

OpenStudy (mathmale):

Most algebra problems involve "solving for an unknown quantity."

OpenStudy (shealyn):

i still got 12.5

OpenStudy (shealyn):

but its not an option

OpenStudy (mathmale):

Would you please share your work? I can tell little or nothing if you only say, "I still got 12.5." How did you get 12.5?

OpenStudy (mathmale):

\[2a^2=625\] must be solved for a^2 first, and then for a. Divide both sides of the equation by 2, to obtain an expression for a^2.

OpenStudy (shealyn):